OpenStudy (anonymous):

Factor and leave the solution in a form such as this: (x+2)^2. x^2-8x+16

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

Could you be more specific?

OpenStudy (anonymous):

Factor and leave the solution in a form such as this: (x+2)^2. ---Factor x^2-8x+16

OpenStudy (anonymous):

(x-4)^2=(x-4)(x-4)=x^2-8x+16. Hope this helps! //Ali